OpenStudy (anonymous):

medal(:

OpenStudy (anonymous):

@pkjag Write an equation for each line in slope-intercept form. 21. slope= ½ and the y-intercept is (0, 6) help with a couple more questions? then you can ask in my inbox

TheSmartOne (thesmartone):

slope intercept form \[(y-y _{1})=m(x-x _{1})\] where m is slope and the point is \[(x _{1},y _{1})\]

OpenStudy (anonymous):

slope= 1/2 sorry don't know why my laptop did that

TheSmartOne (thesmartone):

so \[(y-6)=\frac{ 1 }{ 2 }(x-0)\]

OpenStudy (anonymous):

okay hold on :) ill write it out

OpenStudy (anonymous):

@TheSmartOne so that's the equation in slope-intercept form?

TheSmartOne (thesmartone):

no whoops typo that is the point slope form.

TheSmartOne (thesmartone):

So now solve it. and arrange it into the format of y=mx+b

OpenStudy (anonymous):

oh okay hold on

TheSmartOne (thesmartone):

It is \[y=\frac{ 1 }{2 }x+6\]

OpenStudy (anonymous):

y=1/2x+6? for the y-intercept I have two points (0, 6) do I just put 6?

TheSmartOne (thesmartone):

What do you mean y=1/2x+6 is in slope intercept form...

OpenStudy (anonymous):

(0,6)isn't two points its actually the value 6 on the y-axis, which is the y-intercept, but the question only requires you write the equation

TheSmartOne (thesmartone):

true that ^^^

OpenStudy (anonymous):

oh >_< sorry

OpenStudy (anonymous):

sorry my computer froze I was trying to say I got confused because on my teachers notes she let me use to study the questions I got wrong she wrote "two points"